2. An apple juice producer buys all his apples from a conglomerate of apple growers in one northwest state. The amount of juice squeezed from each of these apples is approximately normally distributed with a mean of 2.15 ounces and a standard deviation of 0.15 ounce.

c. What is the probability that a randomly selected apple will contain more than 2.40 ounces?
Answer #1

2)

Solution :

Given that ,

mean = = 2.15

standard deviation = = 0.15

P(x > 2.40) = 1 - P(x < 2.40)

= 1 - P((x - ) / < (2.40 - 2.15) / 0.15)

= 1 - P(z < 1)

= 1 - 0.8413

= 0.1587

Probability = 0.1587